What effective interest rate per two-months, compounded continuously, is equivalent to a nominal rate of 10% per year? 


The effective interest rate per two-months is _______ %
Answer #1

The formula for Effective interest rate when continuously compounding = ert - 1, where

r: rate per period = 10%/6= 0.01667 and

t: Number of period per year = 6

Effective interest rate with continuous compounding = e(0.01667 x 6) - 1 = e0.10002 - 1 = 0.10519 = 10.52%

Please don't forget to rate the answer if its helpful, thank you.